Circular No. 8652

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 71P/CLARK This comet has been recovered at total mag ∼ 17.5–18 via observations

at several observatories during Dec. 22–30, with the astrometry, the following orbital elements by B. G. Marsden (from 78 observations, 2000–2005; mean residual 0′′.7; nongravitational parameters A_{1} = +1.05 ± 0.19, A_{2} = −0.7600 ± 0.1056), and an ephemeris appearing on MPEC 2005-Y59. The correction to the prediction by S. Nakano (MPC 48382; 2006 Comet Handbook) is ΔT = −0.42 day.

Epoch = 2006 May 25.0 TT

T = 2006 June 6.79246 TT  ω = 208.^{o}74552
e = 0.4997464  Ω = 59.65809 }2000.0
q = 1.5620858 AU  i = 9.48790

a = 3.1225878 AU n^{o} = 0.17862083 P = 5.52 years

COMET C/2005 Y2 (McNAUGHT) R. H. McNaught reports his discovery of a diffuse comet on CCD survey

exposures obtained with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below). Better exposures taken on Dec. 31.44–31.46 UT confirm a coma diameter of ∼ 10′′.

2005 UT  α_{2000}  δ_{2000}  Mag.
Dec. 30.45574  23^{h}07^{m}04.^{s}58  −24^{o}09′24′′.7  17.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-A01.

T = 2006 July 25.690 TT  ω = 295.^{o}386
Ω = 93.331 }2000.0
q = 3.72356 AU  i = 19.076

SUPERNOVAE Several recently discovered supernovae have been found to be of type

II: 2005kk (details on CBET 312), 2005kx (CBET 311), 2005lx (CBET 342), 2005mb (CBET 345), 2005me (CBET 345), 2005mg (CBET 342). Type-Ia supernovae: 2005iq (CBET 278), 2005kc (CBET 286), 2005ls (CBET 324), 2005lt (CBET 319, 321), 2005lu (CBET 321), 2005ms (CBET 345). Type-Ic supernovae: 2005kf (CBET 301), 2005lr (CBET 321).

2005 December 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green